# Improving the Diffusability of Autoencoders

### v2

@@ Abstract
Latent diffusion models have emerged as the leading approach for generating high-quality images and videos, utilizing compressed latent representations to reduce the computational burden of the diffusion process. While recent advancements have primarily focused on scaling diffusion backbones and improving autoencoder reconstruction quality, the interaction between these components has received comparatively less attention. In this work, we perform a spectral analysis of modern autoencoders and identify inordinate high-frequency components in their latent spaces, which are especially pronounced in the autoencoders with a large bottleneck channel size. We hypothesize that this high-frequency component interferes with the coarse-to-fine nature of the diffusion synthesis process and hinders the generation quality. To mitigate the issue, we propose scale equivariance: a simple regularization strategy that aligns latent and RGB spaces across frequencies by enforcing scale equivariance in the decoder. It requires minimal code changes and only up to K autoencoder fine-tuning steps, yet significantly improves generation quality, reducing FID by 19% for image generation on ImageNet-1K and FVD by at least for video generation on Kinetics-700 .
@@ 1 Introduction
Recently, the research community has begun focusing more on improving the autoencoders, recognizing their crucial impact on overall performance, but most effort has been concentrated on enhancing reconstruction quality (Black Forest Labs, 2023; Hong et al., 2022; Agarwal et al., 2025; HaCohen et al., 2024; Chen et al., 2025) and achieving higher compression ratios (Agarwal et al., 2025; HaCohen et al., 2024; Chen et al., 2025) to accelerate the diffusion process. However, we argue that a critical yet under-explored aspect, which we refer to as diffusability 111Diffusability describes how easily a distribution can be modeled by a diffusion process: high diffusability indicates that the distribution is easy to fit, whereas low diffusability makes the process more complex., also plays a key role in determining the utility of autoencoders. Indeed, all three factors—reconstruction quality, compression efficiency, and diffusability—are essential for the practical effectiveness of LDMs. Specifically, inaccurate reconstruction sets an upper bound on generation fidelity, low compression efficiency leads to slow and costly generation, and poor diffusability necessitates the use of heavier, more expensive, and sophisticated diffusion backbones, further limiting LDM quality.
@@ 2 Related Work
Image and video autoencoders. Due to the success of LDMs, a lot of effort has been devoted to the development of better AEs. Image LDMs (Rombach et al., 2022) and early video diffusion models (Blattmann et al., 2023; Guo et al., 2023) employ a spatial AE with a compression ratio of . The rapid advancement of video diffusion models poses the demand for 3D AEs that jointly compress spatial and temporal dimensions to further improve efficiency (Hong et al., 2022; Zhou et al., 2024; Kong et al., 2024). Among them, Open-Sora (Zheng et al., 2024) inherits the spatial AE and trains a decoupled temporal AE on top of its latent space, while others tend to build a hierarchical spatio-temporal AE with 3D causal convolutions (Xing et al., 2024; Wu et al., 2024; Chen et al., 2024a; Zhao et al., 2024; Sadat et al., 2024; Hansen-Estruch et al., 2025). To improve the reconstruction quality, Open-Sora-Plan (Lin et al., 2024) and CosmosTokenizer (Agarwal et al., 2025) propose to employ wavelet transforms to enhance high-frequency details. Another popular trend is to further increase the compression ratio to reduce the number of tokens in the latent space (Xie et al., 2024; Tian et al., 2024; HaCohen et al., 2024), thus enabling a more efficient denoising process. In addition to the continuous AEs explored in this work, multiple discrete AEs (Wang et al., 2024; Tang et al., 2024; Agarwal et al., 2025) are proposed to aid autoregressive tasks. Esteves et al. (2024) leverages a wavelet transform to produce latents corresponding to different frequency components.
@@
Concurrent works. Independently from us, EQ-VAE (Kouzelis et al., 2025) proposes scale equivariance regularization for autoencoders, but with a different motivation of improving their equivariance to spatial transforms.
@@ 3.1 Background: Blockwise 2D DCT
By analyzing RGB images and latents in the DCT frequency domain, we produce a frequency profile that relates to the energy of the signal at every frequency. A zigzag frequency index is used to map each DCT block into a one-dimensional sequence following the standard zigzag ordering as in JPEG (Wallace, 1991), which indexes the DCT coefficients from lowest frequency to highest frequency . Formally, let denote the ranks of the coefficient in ascending frequency order. Given a block, we compute its DCT and produce normalized amplitudes for each frequency component as:
@@ 3.2 Spectral Analysis of the Latent Space
We begin our analysis by observing the frequency profile of the latent space in the Flux (Black Forest Labs, 2023) family of autoencoders to establish a relationship with diffusability. For the purpose of this study, we train a family of FluxAE models with various channel sizes for 100k steps (where performance saturates in this setting) and, for each of them, compute the averaged frequency profile over 256 samples, all channels, and all DCT blocks. Figure 2 presents the frequency profiles of both Flux autoencoders and RGB space, from which we observe: (i) The Flux profile exhibits significantly larger high-frequency components compared to the RGB profile. (ii) As the number of channels in the autoencoder’s bottleneck increases, high-frequency components become more pronounced. This observation is of particular interest as the number of channels is positively correlated with autoencoder’s reconstruction quality.
@@
A common approach to regularizing the latent space in latent diffusion models (LDMs) is to employ a variational autoencoder (VAE) (Kingma & Welling, 2013) framework with a KL divergence term, encouraging the latent distribution to align with a Gaussian prior. This regularization can help to bring the latent distribution closer to the standard Gaussian (Vahdat et al., 2021), simplifying the job for the diffusion process, as the reverse process starts with the same distribution. However, as we show in Figure 3 which compares FluxAE models trained with different levels of KL regularization, higher KL regularization introduces more high frequencies.
@@
As described in (Ning et al., 2024; Dieleman, 2024; Rissanen et al., 2023), diffusion models can be interpreted in the spectral domain as autoregressive processes: when noise level is high, low frequencies are generated, then, as the level of noise lowers during sampling, progressively higher frequencies are generated. This property is desirable, as it allows the model to leverage the cleaner lower frequencies as a conditioning signal for the current prediction. However, the strength of this autoregressive pattern is directly related to the shape of the frequency profile for the signal to generate. Since the white noise that is applied as part of the diffusion process has a flat frequency profile, it follows that the flatter the frequency profile of the signal, the lower the cleanliness of low frequencies that can act as conditioning for the model. For a flat frequency profile, no autoregressive generation is possible as all frequencies would be erased at the same speed by white noise. We also hypothesize that higher frequencies components are harder to model than lower frequency components for the following reasons and thus should be avoided: (i) they have higher dimensionality; (ii) they are generated only in the final steps of sampling, thus must emerge more rapidly; (iii) they are more susceptible to error accumulation over time.
@@ 3.3 Scale Equivariance Regularization
Effective regularization should achieve two key objectives: (i) to suppress high-frequency components in the latent space and (ii) to prevent the decoder from amplifying these components, as their impact on the final result is what ultimately matters. This can be accomplished by aligning the spectral properties of the latent and RGB spaces at different frequencies. A way to achieve this consists in explicitly chopping off a portion of the high frequencies in both spaces and training the decoder to reconstruct the truncated RGB signal from the truncated latent representation. Our preliminary experiments demonstrated that an autoencoder can easily learn to alter its latent frequency profile to encode the inputs in the low-frequency region of the spectrum without sacrificing the reconstruction quality much (see Figure 5). While this regularization, which we name Chopping High Frequencies (CHF), improves the spectrum, we develop a much simpler procedure to achieve the same effect without the need to perform the error-prone DCT transform (the details of CHF are described in Appendix B). The simplest way to achieve high-frequency truncation is through direct downsampling, which we discuss next.
@@ 4.1 Improving Existing Autoencoders
For the LDM benchmark, we utilized ImageNet (Deng et al., 2009) at resolution. We used the DiT (Peebles & Xie, 2022) model as the backbone since it is the most popular modern latent diffusion backbone. Compared to the original paper, we incorporated several recent advancements into the DiT architecture to improve the baseline performance, as described in Appendix A. Qualitative samples from DiT-XL/2 are provided in Figure 7. The results are shown in Table 1. One can observe that our proposed regularization greatly improves the diffusability of the downstream LDM model, allowing to achieve 19% lower FID compared to the vanilla Flux AE and 8% lower FID compared to the Flux AE, fine-tuned in our training pipeline without the SE regularization.
@@
Improving video autoencoders. For video autoencoders, we used CogVideoX-AE (Hong et al., 2022) (CV-AE) with compression and 16 latent channels and LTX-AE (HaCohen et al., 2024) with compression and 32 latent channels. The latter serves as a strong high-compression autoencoder baseline. For all the experiments, we fine-tune them for 20,000 training steps on the joint image and video dataset with the batch size of 32. Image batches are treated as single-frame videos which is possible due to the causal structure of the video autoencoders (Yu et al., 2023).
@@ 5 Conclusion
Appendix A Implementation Details
@@
For CV-AE, since it is considerably slower than other autoencoders, we trained LDMs on pre-extracted latents. For this, we pre-extracted them on random 17-frames clips. In essence, this reduces
@@
Appendix B Additional Exploration
@@
B.1 Explicitly Chopping off High-Frequency Components
@@
B.2 Soft Penalty for High-Frequency Components
@@
To ablate its importance, we trained DiT-B/1 model on top of FluxAE models, fine-tuned with a different strength . The results are presented in Table 7.
@@ B.3 ImageNet experiments
B.4 Ablating regularization strength
@@
Appendix C Additional visualizations
This section provides additional visualizations for the LDM experiments.
Appendix D Limitations.
1.
2.
3.
4.
We leave the exploration of these limitations for future work.

### v3

@@ Abstract
Latent diffusion models have emerged as the leading approach for generating high-quality images and videos, utilizing compressed latent representations to reduce the computational burden of the diffusion process. While recent advancements have primarily focused on scaling diffusion backbones and improving autoencoder reconstruction quality, the interaction between these components has received comparatively less attention. In this work, we perform a spectral analysis of modern autoencoders and identify inordinate high-frequency components in their latent spaces, which are especially pronounced in the autoencoders with a large bottleneck channel size. We hypothesize that this high-frequency component interferes with the coarse-to-fine nature of the diffusion synthesis process and hinders the generation quality. To mitigate the issue, we propose scale equivariance: a simple regularization strategy that aligns latent and RGB spaces across frequencies by enforcing scale equivariance in the decoder. It requires minimal code changes and only up to K autoencoder fine-tuning steps, yet significantly improves generation quality, reducing FID by 19% for image generation on ImageNet-1K and FVD by at least for video generation on Kinetics-700 . The source code is available at https://github.com/snap-research/diffusability.
@@ 1 Introduction
Recently, the research community has begun focusing more on improving the autoencoders, recognizing their crucial impact on overall performance, but most effort has been concentrated on enhancing reconstruction quality (Black Forest Labs, 2023; Hong et al., 2022; Agarwal et al., 2025; HaCohen et al., 2024; Chen et al., 2025) and achieving higher compression ratios (Agarwal et al., 2025; HaCohen et al., 2024; Chen et al., 2025) to accelerate the diffusion process. However, we argue that a critical yet under-explored aspect, which we refer to as diffusability 111Diffusability describes how easily a distribution can be modeled by diffusion: high diffusability indicates that the distribution is easy to fit, while low diffusability implies a higher difficulty., also plays a key role in determining the utility of autoencoders. Indeed, all three factors—reconstruction quality, compression efficiency, and diffusability—are essential for the practical effectiveness of LDMs. Specifically, inaccurate reconstruction sets an upper bound on generation fidelity, low compression efficiency leads to slow and costly generation, and poor diffusability necessitates the use of heavier, more expensive, and sophisticated diffusion backbones, further limiting LDM quality.
@@ 2 Related Work
Image and video autoencoders. Due to the success of LDMs, a lot of effort has been devoted to the development of better AEs. Image LDMs (Rombach et al., 2022) and early video diffusion models (Blattmann et al., 2023; Guo et al., 2023) employ a spatial AE with a compression ratio of . The rapid advancement of video diffusion models poses the demand for 3D AEs that jointly compress spatial and temporal dimensions to further improve efficiency (Hong et al., 2022; Zhou et al., 2024; Kong et al., 2024). Among them, Open-Sora (Zheng et al., 2024) inherits the spatial AE and trains a decoupled temporal AE on top of its latent space, while others tend to build a hierarchical spatio-temporal AE with 3D causal convolutions (Xing et al., 2024; Wu et al., 2024; Chen et al., 2024a; Zhao et al., 2024; Sadat et al., 2024; Hansen-Estruch et al., 2025). To accelerate the model and improve the reconstruction, Open-Sora-Plan (Lin et al., 2024) and CosmosTokenizer (Agarwal et al., 2025) propose to employ the wavelet transform of the input. Another popular trend is to further increase the compression ratio to reduce the number of tokens in the latent space (Xie et al., 2024; Tian et al., 2024; HaCohen et al., 2024), thus enabling a more efficient denoising process. In addition to the continuous AEs explored in this work, multiple discrete AEs (Wang et al., 2024; Tang et al., 2024; Agarwal et al., 2025) are proposed to aid autoregressive tasks. Esteves et al. (2024) leverages a wavelet transform to produce latents corresponding to different frequency components.
@@
Concurrent works. Independently from us, AF-LDM (Zhou et al., 2025) enforces shift equivariance in both the autoencoder and LDM, and EQ-VAE (Kouzelis et al., 2025) proposes scale/shift equivariance regularization for autoencoders, but with a different motivation of improving the models’ equivariance to spatial transformations.
@@ 3.1 Background: Blockwise 2D DCT
By analyzing RGB images and latents in the DCT frequency domain, we produce a frequency profile that relates to the energy of the signal at every frequency. A zigzag frequency index is used to map each DCT block into a one-dimensional sequence following the standard zigzag ordering as in JPEG (Wallace, 1991), which indexes the DCT coefficients from lowest frequency to highest one . Formally, let denote the ranks of the coefficients in the ascending frequency order. Given a block, we compute its DCT and produce the normalized amplitudes for each frequency component as:
@@ 3.2 Spectral Analysis of the Latent Space
We begin our analysis by observing the frequency profile of the latent space in the Flux (Black Forest Labs, 2023) family of autoencoders to establish a relationship with diffusability. For the purpose of this study, we train a family of FluxAE models with various channel sizes for 100k steps (where performance saturates in this setting) and, for each of them, compute the averaged frequency profile over 256 samples, all channels, and all DCT blocks. Figure 2 presents the frequency profiles of both Flux autoencoders and RGB space, from which we observe: (i) The Flux profile exhibits significantly larger high-frequency components compared to the RGB profile. (ii) As the number of channels in the autoencoder’s bottleneck increases, high-frequency components become more pronounced. We hypothesize that a larger bottleneck allows the autoencoder to capture finer, high-frequency details. Initially, limited capacity prioritizes smoother, low-frequency information. But as the capacity expands, the model encodes additional high-frequency content, distributing it unevenly and in the unstructured manner across channels. This finding is of particular interest as the number of channels is positively correlated with autoencoder’s reconstruction quality.
@@
A common way to regularize the latent space for latent diffusion models (LDMs) is to employ a variational autoencoder (VAE) (Kingma & Welling, 2013) framework with a KL divergence term, encouraging the latent distribution to align with the standard Gaussian prior. Since the reverse diffusion process also starts with the same standard normal distribution, such KL regularization is argued to simplify the job for the diffusion model (Vahdat et al., 2021), since it now has less “work” to do. However, as we show in Figure 3 which compares FluxAE with varying levels of KL regularization, higher KL regularization introduces more high frequencies due to the underlying noise addition process, creating a harmful side-effect, reducing diffusability.
@@
Previous work (Ning et al., 2024; Dieleman, 2024; Rissanen et al., 2023) interprets diffusion models as autoregressive ones, but in the spectral domain: when noise level is high, low frequencies are generated, then, as the level of noise lowers during sampling, progressively higher frequencies are generated. This is an attractive property since it allows the model to leverage the cleaner lower frequencies as a conditioning signal for the current prediction. However, the strength of this autoregressive pattern is directly related to the shape of the frequency profile for the signal to generate. Since the white noise that is applied as part of the diffusion process has a flat frequency profile, it follows that the flatter the frequency profile of the signal, the lower the cleanliness of low frequencies that can act as conditioning for the model. For a flat frequency profile, no autoregressive generation is possible as all frequencies would be erased at the same speed by white noise. We also hypothesize that higher frequencies components are harder to model than lower frequency components for the following reasons and thus should be avoided: (i) they have higher dimensionality222E.g., going from to resolution adds little structural content, but increases the dimensionality times.; (ii) they are generated only in the final steps of sampling, thus must emerge more rapidly; (iii) they are more susceptible to error accumulation over time (Li & van der Schaar, 2024).
@@ 3.3 Scale Equivariance Regularization
Effective regularization should achieve two key objectives: (i) to suppress high-frequency components in the latent space and (ii) to prevent the decoder from amplifying these components, as their impact on the final result is what ultimately matters. This can be accomplished by aligning the spectral properties of the latent and RGB spaces at different frequencies. A way to achieve this consists in explicitly chopping off a portion of the high frequencies in both spaces and training the decoder to reconstruct the truncated RGB signal from the truncated latent representation. Our preliminary experiments demonstrated that an autoencoder can easily learn to alter its latent frequency profile to encode the inputs in the low-frequency region of the spectrum without sacrificing the reconstruction quality much (see Figure 5). While this regularization, which we name Chopping High Frequencies (CHF), improves the spectrum, we develop a much simpler procedure to achieve the same effect without the need to perform the error-prone DCT transform (the details of CHF are described in Appendix C). The simplest way to achieve high-frequency truncation is through direct downsampling, which we discuss next.
@@ 4.1 Improving Existing Autoencoders
For the LDM benchmark, we utilized ImageNet (Deng et al., 2009) at resolution. We used the DiT (Peebles & Xie, 2022) model as the backbone since it is the most popular modern latent diffusion backbone. Compared to the original paper, we incorporated several recent advancements into the DiT architecture to improve the baseline performance, as described in Appendix B. Qualitative samples from DiT-XL/2 are provided in Figure 7. The results are shown in Table 1. One can observe that our proposed regularization greatly improves the diffusability of the downstream LDM model, allowing to achieve 19% lower FID compared to the vanilla Flux AE and 8% lower FID compared to the Flux AE, fine-tuned in our training pipeline without the SE regularization.
@@
Improving video autoencoders. For video autoencoders, we used CogVideoX-AE (Hong et al., 2022) (CV-AE) with compression and 16 latent channels and LTX-AE (HaCohen et al., 2024) with compression and 32 latent channels. The latter serves as a strong high-compression autoencoder baseline. All the video AEs are fine-tuned for 20,000 training steps on the joint image and video dataset with the batch size of 32. Image batches are treated as single-frame videos which is possible due to the causal structure of the video autoencoders (Yu et al., 2023).
@@ 4.2 Ablations
Can the quality improvement come from implicit time shifting? SE smoothes the latent space, which can result in “implicit time shifting” (Gao et al., 2024): it eliminates high-frequency components from the latents, allowing the model to spend more compute on low-frequency ones, especially at inference time (see B). We trained a family of DiT-B/2 models for various location and scale coefficients for logit-normal noise distribution (Esser et al., 2024). Figure 10 presents the results with 128 inference steps (and without CFG) for time shifting coefficients of . Our regularized FluxAE-FT+SE achieves the best results across all the setups, confirming that our improved quality is not due to improved diffusability, rather than implicit time shifting. Note that our original noise schedule was tuned for the vanilla setup, which is why the default location of 0 and scale of 1 perform the best for non-regularized AEs.
@@ 5 Conclusion
Appendix A Limitations.
1.
2.
3.
4.
We leave the exploration of these limitations for future work.
Appendix B Implementation Details
@@
Following (Jabri et al., 2023; Chen & Li, 2023; Gupta et al., 2024; Menapace et al., 2024), we employ self-conditioning for our DiT training and inference. During training with a 90% probability, we run an auxiliary forward pass with the DiT model, take its activations from the last block (i.e., right before the “unpatchify” projection), project them with a linear layer and add as residuals to the input tokens after patchification of the main training forward pass. For that auxiliary forward pass, following RIN (Jabri et al., 2023), we use the same noise level and “no-grad” context (i.e., we do not backpropagate through the auxiliary forward pass).
@@
For CV-AE, since it is considerably slower than other autoencoders, we trained LDMs on pre-extracted latents. For this, we pre-extracted them on random 17-frames clips. In essence, this reduces the total dataset size, but since we do the same procedure for the entire CogVideoX-AE family, the models are comparable between each other.
@@
Time shifting for rectified flow. In Section 4.2, we discussed the influence of time shifting on the DiT model. Here, we provided the details of its implementation. We adapt the time shifting schedule from Lumina-T2X (Gao et al., 2024) for rectified flow parametrization by rescaling the time steps at inference time using the following formula:
where is the time step, going from 1 (full noise) to 0 (clean image), and is the shifting coefficient. One can see that for , , which is the default unshifted behaviour.
Appendix C Additional Experiments
@@
C.1 Explicitly Chopping off High-Frequency Components
@@
C.2 Soft Penalty for High-Frequency Components
@@
To ablate its importance, we trained DiT-B/2 model on top of FluxAE models, fine-tuned with a different strength . The results are presented in Table 7.
@@ C.3 ImageNet experiments
C.4 Ablating regularization strength
@@
C.5 From scratch training
To conduct whether our regularization provides benefits for from scratch training and when the same data is being used for both AE and LDM, we conducted additional experiments for FluxAE and CogVideoX-AE. We trained FluxAE from scratch on ImageNet with/without our regularization and then trained DiT-B/2 on top of the latent spaces of these two autoencoders. Then, we did the same for CogVideoX-AE for Kinetics-700 . The results of AE reconstruction performance (after 300K training steps for FluxAE and 60K training steps for CogVideoX-AE) and LDM generation performance after 400K training steps are presented in Table 10. One can observe that our regularization improves the performance in both cases.
| Method | PSNR | LPIPS | g | g | g |
| FluxAE | 29.39 | 0.059 | 502.44 | 25.43 | |
| +FT-SE (ours) | 28.73 | 0.065 | 472.26 | 22.32 | |
| CogVideoX-AE | 37.08 | 0.049 | 614.95 | 19.87 | 336.33 |
| +FT-SE (ours) | 36.61 | 0.053 | 515.68 | 17.40 | 237.28 |
C.6 Compute cost analysis
One might argue that the improved performance of our regularization comes from the extra computation which the autoencoder is using during training. In this section, we show that it is not the case by providing an analysis of its computational cost.
We measured the FLOPs of FluxAE (for the batch size of 1 and resolution of 2562) using the popular fvcore library. The entire encoder-decoder pass has 447 GFLOPS, and is split between the encoder/decoder as 136 vs 311 GFLOPS. Our regularization reuses the encoder pass and only runs the decoder with or reduced resolution (the scale sampled randomly during training). This results in 77.6 or 19.4 extra GFLOPs of the decoder, which is almost exactly 1/4 or 1/16 of the decoder compute or +17% or 4.5% of the total forward pass. Since we sample or downsampling factor equally randomly, this results in 10.75% of total FLOPs overhead for our regularization.
To strengthen our point even further, we ran an experiment where the baseline FluxAE was fine-tuned strictly for 2 times longer (for K instead of K iterations). The resulted DiT-B/2 model achieved and of only 33.99 and 642.7 versus the scores of 25.87 and 551.27 respectively for our regularized FluxAE+FT-SE model, fine-tuned for only 10K steps.
Appendix D Additional visualizations
D.1 Spectra visualizations
D.2 Samples visualizations
Appendix E Failed experiments
Over the course of the project, we explored several other ideas to regularize the latent space of autoencoders. While they did not pan out that well, we still discuss them in this section to spur potential future exploration.
MiniLDM-regulzation training. LSGM (Vahdat et al., 2021) showed that the correct objective of LDM training optimizes both autoencoder and LDM at the same time:
where is the ELBO objective, are encoder/decoder, and is the LDM (or any other latent generative model). This motivated us to explore training a small LDM model together with the autoencoder expecting that it would make its latent space more diffusable. However, even with extensive hyperparameter search, it was not outperforming the vanilla baseline.
Lipszhitz regularization. Dao et al. (2023) derived the following upper bound on the Wasserstein distance between the ground truth and recovered distributions, when the distribution is learned in the latent space of an autoencoder with encoder and decoder :
where and are the ground-truth and recovered velocity estimators in the rectified flow framework (Liu et al., 2022), is the Lipschitz constant of the decoder and is the Lipschitz constant of the learned velocity estimator . This upper bound inspired us to minimize the Lipschitz constant of the decoder. To do this, we used -regularization (Mescheder et al., 2018) from the GAN literature (Goodfellow et al., 2014). This was yielding promising initial results and working almost on par with the scale equivariance regularization, but second-order differentiation for regularization was entailing much slower training speed and engineering struggles (it has poor compatibility with FSDP). This is why we proceeded with scale equivariance which was performing slightly better and much simpler conceptually.
Temporal scale equivariance. We briefly explored temporal scale-equivariance regularization, but surprisingly, it was not leading to improved results. We hypothesize that the temporal domain is of different nature compared to the spatial one since the temporal high-frequency components are more noticeable by human eye than the low-frequency ones.